# RevealNet: Distributed Traffic Correlation for Attack Attribution on Programmable Networks

### v1

@@ Abstract
Network attackers have increasingly resorted to proxy chains, VPNs, and anonymity networks to conceal their activities. To tackle this issue, past research has explored the applicability of traffic correlation techniques to perform attack attribution, i.e., to identify an attacker’s true network location. However, current traffic correlation approaches rely on well-provisioned and centralized systems that ingest flows from multiple network probes to compute correlation scores. Unfortunately, this makes correlation efforts scale poorly for large high-speed networks.
@@
In this paper, we propose RevealNet, a decentralized framework for attack attribution that orchestrates a fleet of P4-programmable switches to perform traffic correlation. RevealNet builds on a set of correlation primitives inspired by prior work on computing and comparing flow sketches—compact summaries of flows’ key characteristics—to enable efficient, distributed, in-network traffic correlation. Our evaluation suggests that RevealNet achieves comparable accuracy to centralized attack attribution systems while significantly reducing both the computational complexity and bandwidth overheads imposed by correlation tasks.
Keywords:
Programmable switches Sketches Traffic correlation.1 Introduction
@@
In recent years, network attackers have increasingly relied on proxies [26], VPNs [5], and anonymity networks [27, 18], to conceal their identities while engaging in malicious network activities. These relay-based anonymization tools route traffic through multiple intermediary servers, thereby obscuring an attacker’s original IP address (i.e., a so-called stepping-stone attack [52]). Consequently, traditional approaches to identify the source of an attack, such as analyzing a flow’s 5-tuple data, fail to trace malicious traffic effectively. This makes it challenging for network operators to perform attack attribution (i.e., to locate the true source of attacks), thereby preventing coordinated response efforts (e.g., via information sharing between ISPs), legal action, or better insights into attackers’ tactics [11].
@@
To uncover the sources behind malicious and anonymized traffic [43, 13], researchers have increasingly relied on traffic correlation techniques. These techniques aim to deanonymize malicious sources of traffic by analyzing and matching their traffic patterns (such as a flow’s packets’ timing and direction, and/or communication volume [29, 25]) as observed by multiple probe nodes spread across the network [8, 38]. Previous studies developed statistical [28, 22, 36] and machine learning-based methods [29, 32, 9] to improve correlation accuracy. However, these methods require transmitting flows’ features (as observed by the probes) to a central correlator node, responsible for processing such features, thus leading to substantial network bandwidth and computational overheads. While decentralized approaches have been discussed [36, 30], they mostly involve the partitioning of correlation tasks among multiple correlator nodes, thus still requiring probes to exchange flow features in bulk towards special-purpose servers, hence only partially mitigating scalability concerns (in particular, that of computation).
@@
Addressing the scalability issues of existing attack attribution frameworks has proven particularly challenging in high-speed and large-traffic volume infrastructures (e.g., software-defined and programmable networks, such as those found in 5G deployments), where the volume of telemetry data grows rapidly with link speeds and which require rapid processing capabilities to uphold performance standards [31, 23, 4]. To address constraints on data storage and the bandwidth overheads imposed by telemetry data offloading [50] in the context of attack attribution, researchers have investigated the use of feature aggregation [25] and compression techniques that produce flow sketches [12, 30], i.e., compact representations of flows’ characteristics which can be used for correlation. While a significant step forward, we argue that sketches alone do not fully address the fundamental scalability limitations of attack attribution workloads.
@@
This paper introduces RevealNet, a framework for attack attribution that operates via the decentralized correlation of attacking flows. RevealNet eschews the need for special-purpose correlation nodes and minimizes data exchanges during correlation tasks. At the core of our approach is the realization that, while flow correlation capabilities remain largely unexplored in P4-programmable switches (e.g., Intel Tofino [19], AMD Pensando [1]), these devices are gaining traction in high-performance networks due to their ability to perform complex network security operations with low computational overhead [16, 53]. This raises the question of whether P4 switches can also leverage efficient correlation-focused flow feature extraction primitives—such as flow sketches—to operate as decentralized probe/correlation nodes, without incurring the additional costs of middlebox infrastructures [37] or of offloading feature processing to dedicated servers [40].
@@
Contributions. The contributions of this paper can be summarized as follows:
@@
2 Background and Related Work
2.1 Traffic Correlation
Traffic correlation techniques can be used to analyze traffic patterns and link together flows which are observed at the entry and exit nodes of proxy chains [20]. While many correlation techniques were introduced with the aim of gauging the privacy provided by anonymity [28, 34, 8] and mix networks [33], others were developed with the specific intent to trace stepping-stone network attackers [41, 49, 12]. Below, we describe two main classes of prominent passive flow correlation techniques: a) those that use fine-grained per-packet data for higher accuracy at the cost of increased storage, and; b) those that rely on coarse-grained per-flow data, which are more storage-efficient but are typically less precise.
Flow correlation with fine-grained information. Most studied traffic correlation techniques rely on fine-grained, per-packet information. Zhu et al. [56] leverage per-packet timing information to compute the average traffic rate of flows at different intervals, while Palmieri [36] used wavelet-based analysis to capture timing, size, and rate variations across flows. Recently, researchers adopted deep learning approaches to improve flow correlation, pushing accuracy over that of statistical methods. DeepCorr [29], DeepCoFFEA [32], and ESPRESSO [9] progressively improve accuracy—DeepCorr uses convolutional neural networks (CNNs) to learn correlation functions, DeepCoFFEA further introduces feature embedding and voting mechanisms, and ESPRESSO combines transformer models and CNNs to capture both global and local traffic patterns.
While the above approaches yield high accuracy, they rely on the collection, communication, and processing of fine-grained information (direction, size, and timing) about packets in a trace, making them costly to deploy at scale (§3,[30]).
@@ 2.1 Traffic Correlation
In Nasr et al. [30] and Lopes et al. [25], however, flows’ succinct representations are only generated after the initial collection of per-packet information. Still, these compact structures may reveal useful for correlation efforts in the scope of stepping-stone detection, should one be able to compute these representations on-the-fly, eschewing the need to store per-packet information. Inspired by these works, we conjecture that these techniques can help reduce memory use at traffic collection nodes and correlate flows using limited flow data (§4.4).
2.2 P4 Switches as a Platform for Traffic Analysis
@@
Primer on P4-programmable switches. P4 switches cleanly separate the responsibilities of the network’s data and control planes. The data plane is optimized for line-rate packet forwarding and allows for programmable, per-packet operations that enable feature extraction without compromising throughput. In turn, the control plane manages rule installation and updates, supporting adaptive responses to changing traffic patterns. P4 switches, such as the Intel Tofino [19] or AMD Pensando [1] devices, move packets through a multi-stage pipeline before forwarding them [7]. Ingress and egress pipelines employ match-action units to handle packet forwarding and programmable logic. After incoming packets are parsed, their headers and metadata can match a given table, whose entry will map to an action unit. Actions can alter packet header fields and modify stateful memory (e.g., increment register counters). Although matching tables and other P4 objects are instantiated inside match-action units, they are populated by the control plane at and throughout run-time.
@@
Despite their benefits, P4 switches bring limitations that restrict programmability, including the lack of dynamic data structures, no support for floating-point arithmetic, limited memory capacity (256MBs SRAM), and tight computational constraints that allow only simple operations per pipeline stage [54, 3]. These constraints pose implementation challenges to P4 programs that require complex flow feature processing and storage, which are typically implemented via clever “hacks” and workarounds [47, 2, 54, 48]. In our proposed design (§4), we leverage similar approaches to make an efficient use of the limited memory and computation primitives in P4 switches to compute and store flow features.
@@
Traffic analysis on P4 switches. P4 switches have been increasingly employed for traffic analysis tasks [10, 55], including traffic classification [25, 45, 46, 54], covert channel detection [47], and DDoS mitigation [24, 44]. Seminal systems in this space focused on extracting fine-grained traffic features within the data plane, and then offloading them to the switches’ control plane to support a range of security-focused tasks. For enhancing traffic analysis capabilities, researchers worked towards implementing efficient data structures and scalable storage management mechanisms for handling many concurrent flows [51], as well as making significant strides for running classifiers in the data plane [54, 48].
@@
To the best of our knowledge, the use of P4 switches has not yet been applied to the problem of flow correlation. The closest work to our setting is DELTA [21], a system where P4-programmable switches are configured to independently identify the establishment of VoIP calls between peers across the network, and then orchestrated to exchange call information to identify the users engaged in communication. However, DELTA does not extract packet or flow-based features that would apply for generic flow correlation tasks. This gap presents an opportunity to study whether P4 switches can act as the backbone of scalable and decentralized infrastructure for attack attribution in high-speed networks (§4).
3 The Attack Attribution Problem
@@
Our scenario models a networked environment such as the Internet, consisting of a set of interconnected networks . Each network is managed by an operator, e.g., an Internet Service Provider (ISP), a cloud service provider (CSP), or a university campus/enterprise network administrator. These networks may deploy local intrusion detection systems (IDSes) in firewalls, servers, or end-hosts, and share telemetry or flow-level metadata to support collaborative cybersecurity operations. We name these cooperating networks.
@@
To perform attack attribution, a network operator extracts traffic features from the malicious flow—denoted —as observed in the attacked network . These features are then correlated with those of other flows originated within cooperating networks in . Here, denotes a flow originated within network , and whose source IP is . Let denote the correlation between the features of and . Correlation may reflect similar traffic timing and volume characteristics, indicating that and may belong to the same communication path or originate from a common source, as observed from different vantage points. Hence, the attack attribution problem can be formalized as identifying the probable attacker’s IP address , as follows:
, where is a tunable flow similarity threshold.
Scalability challenges of centralized attack attribution. Figure 1 shows a centralized flow correlation architecture for attack attribution. The IDS in the attacked network instructs its probe (step ) to send the attacking flow’s feature vector to a central correlator (step ), while cooperating networks forward feature vectors for all of their outgoing flows (step ). The correlator then computes similarity scores between the attacking flow and those from cooperating networks. This design incurs a computation and communication in the order of – the total number of outgoing flows observed by the cooperating networks. As a result, centralized architectures create significant bottlenecks, requiring powerful correlator nodes and high-capacity links. In the next section, we present RevealNet, a distributed framework that addresses these limitations.
4 RevealNet
@@
Next, we start by describing RevealNet’s architecture before outlining its workflow and providing a comprehensive overview of its core design features.
4.1 Architecture Overview
@@
Programmable switches. P4 switches represent a central component of RevealNet, serving a dual role as feature collectors (probe nodes) and correlation engines (correlator nodes). Importantly, these switches may already be deployed at participating networks to function as border routers and perform packet forwarding, making them a readily available platform for in-network processing. RevealNet leverages this existing infrastructure to extract flow-level features in a per-packet fashion, enabling an efficient feature aggregation directly within the data plane (§2.2). Once instructed to initiate correlation, the P4 switches retrieve the set of feature vectors associated with each flow of interest from the data plane and perform the correlation operations internally, on the switch’s CPU (control plane). This ensures that feature extraction and correlation remain tightly coupled and execute within the switch itself, requiring no additional components.
@@
Correlation manager. The correlation manager acts as a lightweight logical entity that orchestrates the distributed correlation workload. We envision that this component may be operated independently by a neutral third party, such as a trusted consortium, an inter-organizational security alliance, or a national cybersecurity center [42]. This component’s role is limited to coordinating the correlation process: it collects metadata and feature vectors of attacking flows from the probe node within an attacked network, and distributes the offending flow’s feature vector to the correlator nodes run by cooperating networks. The actual correlation is performed at those distributed nodes, deliberately avoiding centralized resource-intensive computations. Once the distributed correlation is complete, results are reported back to the correlation manager.
4.2 Operational Workflow
We now outline the operational workflow of our framework. Figure 2 illustrates each of the steps composing RevealNet’s sequential workflow.
Flow features’ extraction. In steady-state, all probe nodes within a cooperating network will produce a compact representation of the features for each flow concurrently crossing the switch at a given point in time (step ). Given programmable switches’ memory limitations, this compact representation, which we refer to as a feature vector, is ephemeral, potentially being replaced in an LRU fashion as flows are terminated [2]. Feature vectors will be used for correlating flows as part of the attack attribution process once an attacking flow is detected within a cooperating network under attack. As we describe later on, RevealNet is compatible with multiple compact representations of a flow’s feature vector (§4.4), enabling us to explore different memory/correlation accuracy trade-offs.
@@ 4.2 Operational Workflow
Propagation of the attacking flows’ features. As requested by the correlation manager, the probe node within the attacked network sends it a feature vector that characterizes the attacking flow (step ).
Correlation directive. Upon receiving the feature vector of the attacking flow, the correlation manager forwards it to the P4 switches (which will now act as correlator nodes) that front-face each cooperating network. In addition, the correlation manager distributes the attacking flow’s details (i.e., the flow’s start times and communication volume) to the same switches, enabling them to preemptively identify which (outgoing) recorded flows have similar start/end times as the offending flow, and reason about heuristic optimizations for the local correlation procedure (§4.5). Finally, the correlation manager instructs these switches to initiate their local correlation process (§4.4) for the attacking flow (step ).
Flow correlation and reporting. Flows with correlation scores matching and/or exceeding a set threshold are flagged as potentially correlated, and any associated 5-tuples are sent back to the correlation manager for concluding the attack attribution process (step ).
4.3 Flow Identification and Tracking
A key operation underlying the execution of RevealNet is that of performing an efficient flow identification within P4 switches, so that accurate per-flow feature vectors can be computed and stored within each switch’s data plane. Still, previous prototypes for ML-based flow analysis schemes for P4 switches side-step the problem of unique flow identification, assuming that some process will be in place to perform this identification (e.g., [2]), or forcing the initial packet of a new connection to move through the control plane for further processing and mapping (e.g., [47]), thus causing delays upon connection establishment in high-speed networks with low latency requirements. Thus, while apparently straightforward, producing a 1-to-1 match between new flows and a memory region that can accommodate for a flow’s compact feature representation is non-trivial should one wish to avoid delays in packet processing and/or feature corruption, e.g., caused by hash-based indexing methods that may lead to collisions [17, 14].
RevealNet’s flow identification pipeline. We now detail how RevealNet tackles flow identification. Figure 3 illustrates the steps of this procedure.

### v2

@@ Abstract
Network attackers have increasingly resorted to proxy chains, VPNs, and anonymity networks to conceal their activities. To tackle this issue, past research has explored the use of traffic correlation techniques to perform attack attribution, i.e., to identify an attacker’s true network location. However, current traffic correlation approaches rely on well-provisioned and centralized systems that ingest flows from multiple network probes to compute correlation scores. Unfortunately, this makes correlation efforts scale poorly for large high-speed networks.
@@
In this paper, we propose RevealNet, a decentralized framework for attack attribution that orchestrates a fleet of P4-programmable switches to perform traffic correlation. We build on top of a set of correlation primitives inspired by prior work on computing and comparing flow sketches—compact summaries of flows’ characteristics—to enable efficient, distributed, in-network traffic correlation. Our evaluation suggests that RevealNet achieves comparable accuracy to centralized attack attribution systems while significantly reducing the computational complexity and bandwidth overheads imposed by correlation tasks.
I Introduction
@@
In recent years, network attackers have increasingly relied on proxies, VPNs, and anonymity networks, to conceal their identities while engaging in malicious network activities. These anonymization tools route traffic through multiple intermediary servers, thereby obscuring an attacker’s original IP address (i.e., a so-called stepping-stone attack [1]). Consequently, traditional approaches to identify the source of an attack, such as analyzing a flow’s 5-tuple data, fail to trace malicious traffic effectively. This makes it challenging for network operators to perform attack attribution (i.e., to locate the true source of attacks), preventing coordinated response efforts (e.g., via information sharing between ISPs), legal action, or better insights into attackers’ tactics [2].
@@
To uncover the sources behind malicious and anonymized traffic [3], researchers have increasingly relied on traffic correlation techniques. These techniques aim to deanonymize malicious sources of traffic by analyzing and matching their traffic patterns (such as a flow’s packets’ timing and direction, and/or communication volume) as observed by multiple probe nodes spread across the network [4, 5]. Previous studies developed statistical [6, 7] and machine learning-based methods [4, 8] to improve correlation accuracy. However, these methods require transmitting flows’ features (as observed by the probes) to a central correlator node, responsible for processing such features, thus leading to substantial network bandwidth and computational overheads. While decentralized approaches have been discussed [7, 9], they mostly involve the partitioning of correlation tasks among multiple correlator nodes and still require probes to exchange flow features in bulk towards special-purpose servers, hence only partially mitigating scalability concerns (in particular, that of computation).
@@
Addressing the scalability issues of existing attack attribution frameworks has proven particularly challenging in high-speed and large-traffic volume infrastructures (e.g., software-defined and programmable networks, such as those found in 5G deployments), where the volume of telemetry data grows rapidly with link speeds and which require rapid processing capabilities to uphold performance standards [10]. To address constraints on data storage and the bandwidth overheads imposed by telemetry data offloading in the context of attack attribution, researchers have investigated the use of feature aggregation [5] and compression techniques that produce flow sketches [11, 9], i.e., compact representations of flows’ characteristics which can be used for correlation. While a significant step forward, we argue that sketches alone do not fully address the fundamental scalability limitations of attack attribution workloads.
@@
This paper introduces RevealNet, a framework for attack attribution that operates via the decentralized correlation of attacking flows. RevealNet eschews the need for special-purpose correlation nodes and minimizes data exchanges during correlation tasks. At the core of our approach is the realization that, while flow correlation capabilities remain largely unexplored in P4-programmable switches (e.g., Intel Tofino, AMD Pensando), these devices are gaining traction in high-performance networks due to their ability to perform complex network security operations with low computational overhead [12]. This raises the question of whether P4 switches can also leverage efficient correlation-focused flow feature extraction primitives—such as flow sketches—to operate as decentralized probe/correlation nodes, without incurring the additional costs of middlebox infrastructures [13] or of offloading feature processing to dedicated servers [14].
@@
Contributions. We summarize our contributions as follows:
@@
II Background and Related Work
II-A Traffic Correlation
Traffic correlation techniques can be used to analyze traffic patterns and link together flows which are observed at the entry and exit nodes of proxy chains. While some correlation schemes helped gauge the privacy provided by anonymity [6] and mix networks [15], others were developed with the intent to trace stepping-stone attackers [16, 11]. Below, we describe two main classes of prominent passive flow correlation techniques: a) those that use fine-grained per-packet data for higher accuracy at the cost of increased storage, and; b) those that rely on coarse-grained per-flow data, which are more storage-efficient but are typically less precise.
Flow correlation with fine-grained information. Most studied traffic correlation techniques rely on fine-grained, per-packet information. Zhu et al. [17] leverage per-packet timing information to compute the average traffic rate of flows at different intervals, while Palmieri [7] used wavelet-based analysis to capture timing, size, and rate variations across flows. Recently, researchers adopted deep learning to improve flow correlation, pushing accuracy over that of statistical methods. DeepCorr [4] and DeepCoFFEA [8] progressively improve accuracy—DeepCorr uses convolutional neural networks (CNNs) to learn correlation functions, and DeepCoFFEA introduces novel feature embedding and voting mechanisms.
While the above approaches yield high accuracy, they rely on the collection, communication, and processing of fine-grained information (direction, size, and timing) about packets in a trace, making them costly to deploy at scale (§III,[9]).
@@ II-A Traffic Correlation
In Nasr et al. [9] and Lopes et al. [5], however, flows’ succinct representations are only generated after the initial collection of per-packet information. Still, these compact structures may reveal useful for correlation efforts in the scope of stepping-stone detection, should one be able to compute these representations on-the-fly, eschewing the need to store per-packet data. Inspired by these works, we conjecture that these techniques can help reduce memory use at traffic collection nodes and correlate flows using limited flow data (§IV-D).
II-B P4 Switches as a Platform for Traffic Analysis
@@
Primer on P4-programmable switches. P4 switches cleanly separate the responsibilities of the network’s data and control planes. The data plane is optimized for line-rate packet forwarding and allows for programmable, per-packet operations that enable feature extraction without compromising throughput. In turn, the control plane manages rule installation and updates, supporting adaptive responses to changing traffic patterns. P4 switches, such as the Intel Tofino or AMD Pensando devices, move packets through a multi-stage pipeline before forwarding them. Ingress and egress pipelines employ match-action units to handle packet forwarding and programmable logic. After incoming packets are parsed, their headers and metadata can match a given table, whose entry will map to an action unit. Actions can alter packet header fields and modify stateful memory (e.g., increment register counters). Although matching tables and other P4 objects are instantiated inside match-action units, they are populated by the control plane at and throughout run-time.
@@
Despite their benefits, P4 switches bring limitations that restrict programmability, including the lack of dynamic data structures, no support for floating-point arithmetic, limited memory capacity (256MBs SRAM), and tight computational constraints that allow only simple operations per pipeline stage. These constraints pose implementation challenges to P4 programs that require complex flow feature processing and storage, which are typically implemented via clever “hacks” and workarounds [19, 20, 21]. In our proposed design (§IV), we leverage similar approaches to make an efficient use of the limited memory and computation primitives in P4 switches to compute and store flow features.
@@
Traffic analysis on P4 switches. P4 switches have been increasingly employed for traffic analysis tasks [22], including traffic classification [20], covert channel detection [19], and DDoS mitigation [23]. Seminal systems in this space focused on extracting fine-grained traffic features within the data plane, and then offloading them to the switches’ control plane to support a range of security-focused tasks. For enhancing traffic analysis capabilities, researchers designed efficient data structures and scalable storage management mechanisms for handling many concurrent flows [24], as well as making significant strides for running classifiers in the data plane [21]. Despite these strides, flow correlation requires maintaining temporal relationships across multiple flows and associating them with their upstream sources, requiring feature aggregation mechanisms that are not addressed in prior work.
@@
To the best of our knowledge, the use of P4 switches has not yet been applied to the problem of flow correlation. The closest work to our setting is DELTA [25], a system where P4-programmable switches are configured to independently identify the establishment of VoIP calls between peers across the network, and then orchestrated to exchange call information to identify the users engaged in communication. However, DELTA does not extract packet or flow-based features that would apply for generic flow correlation tasks. This gap presents an opportunity to study whether P4 switches can act as the backbone of scalable and decentralized infrastructure for attack attribution in high-speed networks (§IV).
III The Attack Attribution Problem
@@
We model a networked environment such as the Internet, consisting of a set of interconnected networks . Each network is managed by an operator, e.g., an Internet Service Provider (ISP), a cloud service provider (CSP), or a university campus/enterprise network administrator. These networks may deploy intrusion detection systems (IDSes) in firewalls, servers, or end-hosts, and share telemetry or flow-level metadata to support collaborative cybersecurity operations. We name these cooperating networks.
@@
To perform attack attribution, a network operator extracts traffic features from the malicious flow—denoted —as observed in the attacked network . These features are then correlated with those of other flows originated within cooperating networks in . Here, denotes a flow originated within network , and whose source IP is . Let denote the correlation between the features of and . Correlation may reflect similar traffic timing and volume characteristics, indicating that and may belong to the same communication path or originate from a common source, as observed from different vantage points. Hence, attack attribution can be formalized as identifying the probable attacker’s IP address , as follows:
, where is a similarity threshold.
Scalability challenges of centralized attack attribution. Figure 1 shows a centralized flow correlation architecture for attack attribution. The IDS in the attacked network instructs its probe (step ) to send the attacking flow’s feature vector to a central correlator (step ), while cooperating networks forward feature vectors for all of their outgoing flows (step ). The correlator then computes similarity scores between the attacking flow and those from cooperating networks. This design incurs a computation and communication in the order of – the total number of outgoing flows observed by the cooperating networks. Thus, centralized architectures create significant bottlenecks, requiring powerful correlator nodes and high-capacity links. In the next section, we present RevealNet, a distributed framework that addresses these limitations.
IV RevealNet
@@
IV-A Architecture Overview
@@
Programmable switches. P4 switches represent a central component of RevealNet, serving a dual role as feature collectors (probe nodes) and correlation engines (correlator nodes). Importantly, these switches may already be deployed at participating networks to function as border routers and perform packet forwarding, making them a readily available platform for in-network processing. RevealNet leverages this existing infrastructure to extract flow-level features in a per-packet fashion, enabling an efficient feature aggregation directly within the data plane (§II-B). Once instructed to initiate correlation, the P4 switches retrieve the set of feature vectors associated with each flow of interest from the data plane and perform the correlation operations internally, on the switch’s CPU (control plane). This ensures that feature extraction and correlation remain tightly coupled and execute within the switch itself, requiring no additional components.
@@
Correlation manager. The correlation manager acts as a lightweight logical entity that orchestrates the distributed correlation workload. We envision that this component may be operated independently by a neutral third party, such as a trusted consortium, an inter-organizational security alliance, or a national cybersecurity center [26] for the purpose of attack attribution, e.g., in the context of international law-enforcement collaborations [2]. This component’s role is limited to coordinating the correlation process: it collects metadata and feature vectors of attacking flows from the probe node within an attacked network, and distributes the offending flow’s feature vector to the correlator nodes run by cooperating networks. The actual correlation is performed at those distributed nodes, deliberately avoiding centralized resource-intensive computations. Once the distributed correlation is complete, results are reported back to the correlation manager.
From a privacy perspective, RevealNet minimizes data disclosure by revealing a potential attack source IP to the correlation manager only when the correlation score between the malicious flow and a flow observed at a cooperating network exceeds a predefined threshold. During correlation, only flow sketches are exchanged, preventing the leakage of personally identifiable information that could be reconstructed from detailed flow information. In addition, RevealNet is compatible with privacy-preserving flow correlation mechanisms, e.g., those based on multi-party computation primitives [27].
IV-B Operational Workflow
We now outline the operational workflow of RevealNet, relying on Figure 2 to illustrate each of these steps.
Flow features’ extraction. In steady-state, all probe nodes within a cooperating network will produce a compact representation of the features for each flow concurrently crossing the switch at a given point in time (step ). Given programmable switches’ memory limitations, this compact representation, which we refer to as a feature vector, is ephemeral, potentially being replaced in an LRU fashion as flows are terminated [20]. Feature vectors will be used for correlating flows as part of the attack attribution process once an attacking flow is detected within a cooperating network under attack. As we describe later on, RevealNet is compatible with multiple compact representations of a flow’s feature vector (§IV-D), enabling us to explore different memory/correlation accuracy trade-offs.
@@ IV-B Operational Workflow
Propagation of the attacking flows’ features. As requested by the correlation manager, the probe node within the attacked network sends the correlation manager a feature vector that characterizes the attacking flow (step ).
Correlation directive. Upon receiving the feature vector of the attacking flow, the correlation manager forwards it to the P4 switches (which will now act as correlator nodes) that front-face each cooperating network. In addition, the correlation manager distributes the attacking flow’s details (i.e., the flow’s start times and communication volume) to the same switches, enabling them to preemptively identify which (outgoing) recorded flows have similar start/end times as the offending flow, and reason about heuristic optimizations for the local correlation procedure (§IV-E). Finally, the correlation manager instructs these switches to initiate their local correlation process (§IV-D) for the attacking flow (step ).
Flow correlation and reporting. Flows with correlation scores matching and/or exceeding a set threshold are flagged as potentially correlated, and any associated 5-tuples are sent back to the correlation manager for concluding the attack attribution process (step ). The manager performs the final attack attribution by aggregating reports from multiple P4 switches (scattered across several networks and sub-networks), identifying candidate attack sources. It could also partially or even fully reconstruct the path of the malicious traffic, even if intermediate cooperating networks did not detect the attack.
A note on insider attacks. Insider attacks represent a simplified case of RevealNet’s workflow. Correlation could be performed within the P4 switching infrastructure of the ISP, assisted by an in-house correlation manager, without the need to share sketches externally. However, correlation accuracy improves with broader visibility, as prior traffic correlation work has shown [5]; restricting correlation to a single network risks missing the true correlated flow if it transits elsewhere. In such cases, weaker local matches may be mistaken for correlations, reducing true positives and increasing false positives.
IV-C Flow Identification and Tracking
A key operation underlying the execution of RevealNet is that of performing an efficient flow identification within P4 switches, so that accurate per-flow feature vectors can be computed and stored within each switch’s data plane. Still, previous prototypes for ML-based flow analysis schemes for P4 switches side-step the problem of unique flow identification, assuming that some process is in place to perform it (e.g., [20]), or forcing the initial packet of a new connection to move through the control plane for further processing and mapping (e.g., [19]), thus causing delays upon connection establishment in high-speed networks with low latency requirements. Thus, while apparently straightforward, producing a 1-to-1 match between new flows and a memory region that can accommodate for a flow’s compact feature representation is non-trivial should one wish to avoid delays in packet processing and/or feature corruption, e.g., caused by hash-based indexing methods that may lead to collisions [28].
RevealNet’s flow identification pipeline. We now detail how RevealNet tackles flow identification, relying on Figure 3.